OpenStudy (anonymous):

College Algebra: How would you find the polynomial equation of the least possible degree with integer coefficients having a greatest common factor of 1 and whose roots include the numbers 0,4,-4?

OpenStudy (mathstudent55):

The polynomial with roots a, b, and c is: (x - a)(x - b)(x - c) = 0

OpenStudy (anonymous):

It would just be (x-0)(x-4)(x+4)?

OpenStudy (anonymous):

And then factored together?

OpenStudy (mathstudent55):

Yes to your first question. Set that equal to zero. Then notice that x - 0 is simply x. Multiply the other two binomials together using FOIL, and then multiply that result by x.

OpenStudy (anonymous):

Okay, I understand now, thank you very much! :)

OpenStudy (anonymous):

x^3-16?

OpenStudy (mathstudent55):

You are welcome.

OpenStudy (mathstudent55):

\(x(x + 4)(x - 4) = 0\) \(x(x^2 - 16) = 0\) \(x^3 - 16x =0\) Don't forget to multiply x by the -16 term also.

OpenStudy (anonymous):

Oh, whoops, thank you again. :)

OpenStudy (mathstudent55):

You're welcome.